OpenStudy (anonymous):

An airplane lands with a speed of 81.9 m/s traveling due south . It comes to rest in 9.49 m. Assuming that the airplane slows with constant acceleration, find the magnitude and the direction of its acceleration.

OpenStudy (anonymous):

\[v_f^2=v_1^2+2ad\] \(v_f=0~m/s\) \(v_i=81.9~m/s\) \(d=9.49~m\) displacement and velocity are south, so acceleration would be as well